# Upper and Lower Bounds on the Smoothed Complexity of the Simplex Method

### v2

@@ 1.3.2 Smoothed Complexity Lower Bound
Our smoothed complexity lower bound (Theorem 49) is based on two geometric observations using the inner and outer radius of the perturbed polytope. For a polytope $P$ and a unit norm ball $\mathbb{B}$, its outer radius with center $x$ is the smallest $R$ such that there exists $P\subseteq R\cdot\mathbb{B}+x$. Its inner radius with center $x$ is the largest $r$ such that $r\cdot\mathbb{B}+x\subseteq P$.
The first observation is that, if a two-dimensional polygon $T$ has inner $\ell_{2}$-radius of $r$ and outer $\ell_{2}$-radius of $(1+\varepsilon)\cdot r$ with respect to the same center, then $T$ has at least $\Omega(\varepsilon^{-1/2})$ edges (Lemma 58). This comes from the fact that every edge of $T$ has length at most $O(r\sqrt{\varepsilon})$, whereas the perimeter of $T$ is at least $2\pi r$.
Second, if two polytopes $Q,\tilde{Q}\subseteq\mathbb{R}^{d}$, each with inner radius $t$, have Hausdorff distance $\varepsilon<t/2$ to each other, then $Q$ will approximate $\tilde{Q}$ in the way that (Lemma 57)
@@
To prove our lower bound, we construct a polytope $Q=\operatorname{conv}(\bar{a}_{1},\ldots,\bar{a}_{n})\subseteq\mathbb{R}^{d}$ and a two-dimensional linear subspace $W$ such that $\Omega(1)\cdot B^{d}_{1}\subseteq Q\subseteq B^{d}_{1}$, and $Q\cap W$ has outer $\ell_{2}$-radius $r>0$ and inner $\ell_{2}$-radius $\frac{r}{(1+4^{-d})}$. Perturbing the vertices of $Q$, we obtain $\tilde{Q}=\operatorname{conv}(a_{1},\ldots,a_{n})$, where $a_{i}\sim N(\bar{a}_{i},\sigma^{2}I_{d\times d})$ for each $i\in[n]$. Note that $Q\subseteq\mathbb{B}_{1}^{d}$ implies that $\bar{a}_{1},\dots,\bar{a}_{n}$ satisfy the normalization requirement in (1). With high probability the Hausdorff distance in $\ell_{1}$ between $Q$ and $\tilde{Q}$ is bounded by $\max_{i\in[n]}\|a_{i}-\tilde{a}_{i}\|_{1}\leq O(\sigma d\sqrt{\log n})$. Using (10), we bound the inner and outer radius of $\tilde{Q}\cap W$. A lower bound on the number of edges of $\tilde{Q}\cap W$ then follows from Lemma 58 as described above.
We remark that the polytopes $Q=\operatorname{conv}(a_{1},\dots,a_{n})\subseteq\mathbb{R}^{d}$ with $n=O(d)$ and two-dimensional subspaces $W$ such that $Q\cap W$ has inner $\ell_{2}$-radius $\frac{r}{1+4^{-d}}$ and outer $\ell_{2}$-radius $r>0$ were first obtained by [BN01] as an extended formulation for a regular $2^{k}$-gon with $O(k)$ variables and $O(k)$ inequalities. Their polytope, however, has an outer and inner radius that differ by a factor $2^{\Omega(k)}$, meaning that we cannot apply Lemma 57 for $\sigma>2^{-k}$. We construct an alternative extended formulation where the ratio between inner and outer $\ell_{1}$-radius is only $O(1)$. With an appropriate scaling to get $Q\subseteq B_{1}^{d}$, we find that the perturbed polytope $\tilde{Q}$ will have intersection $\tilde{Q}\cap W$ with inner radius $\frac{r}{1+4^{-d}}(1-2\varepsilon/t)$ and outer radius $(1+\varepsilon/t)r$, where $\varepsilon=O(\sigma d\sqrt{\log n})$, and thus has $\Omega(\min(\frac{1}{\sqrt{\varepsilon}},2^{d}))$ edges, with high probability.
@@ 2 Preliminaries
We write $[n]:=\{1,\dots,n\}$. The convex hull of vectors $a_{1},\dots,a_{n}$ is denoted $\operatorname{conv}(a_{1},\dots,a_{n})=\operatorname{conv}(a_{i}:i\in[n])$, and similarly the affine hull as $\operatorname{affhull}(a_{i}:i\in[n])$. For sets $A,B\subseteq\mathbb{R}^{d}$, the distance between the two is $\operatorname{dist}(A,B)=\inf_{a\in A,b\in B}\|a-b\|$. For a point $x\in\mathbb{R}^{d}$ we write $\operatorname{dist}(x,A)=\operatorname{dist}(A,x)=\operatorname{dist}(A,\{x\})$.
@@
For a convex body $K\in\mathbb{R}^{d}$, we define $\partial K\subseteq\mathrm{span}(K)$ as the boundary of $K$ in linear subspace spanned by the vectors in $K$.
@@ 2.1 Polytopes
Definition 5 (Polytope).
A convex set $P\subseteq\mathbb{R}^{d}$ is a polyhedron, if it can be expressed as $P=\{x\in\mathbb{R}^{d}:Ax\leq b\}$, for some $A\in\mathbb{R}^{n\times d}$, $b\in\mathbb{R}^{n}$ where $n\in\mathbb{Z}_{+}$. A bounded polyhedron is also called a polytope.
Definition 6 (Valid Condition and Facet).
@@
A subset $F\subseteq P$ is called a face of $P$, if $F=P\cap\{x\in\mathbb{R}^{d}:x^{\top}c=d\}\neq\emptyset$, for some valid condition $x^{\top}c\leq d$. A facet is a $d-1$-dimensional face, a ridge is a $d-2$-dimensional face, an edge is a $1$-dimensional face and a vertex is a 00-dimensional face.
Definition 7 (Polar dual of a convex body).
@@
Fact 8 (Polar dual of polytope).
@@ 2.2 Probability Distributions
Definition 11 (Gaussian distribution).
The $d$-dimensional Gaussian distribution $\mathcal{N}_{d}(\bar{a},\sigma^{2}I)$ with support on $\mathbb{R}^{d}$, mean $\bar{a}\in\mathbb{R}^{d}$, and standard deviation $\sigma$, is defined by the probability density function
@@
Lemma 12 (Gaussian tail bound).
@@
Corollary 13 (Global diameter of Gaussian random variables).
@@
Proof.
From Lemma 12, we have for each $i\in[n]$ that
@@
Then the statement follows from the union bound. ∎
A helpful technical substitute for the Gaussian distribution was introduced by [DH18]:
Definition 14 ($(\sigma,r)$-Laplace-Gaussian distribution).
@@
Definition 15 ($L$-log-Lipschitz random variable).
@@
Lemma 16 (Properties of Laplace-Gaussian random variables, Lemma 45 of [DH18]).
Given any $n\geq d$ , $\sigma>0$. Let $a_{1},\ldots,a_{n}\in\mathbb{R}^{d}$ be independent random variables each sampled form $LG_{d}(\bar{a},\sigma,4\sigma\sqrt{d\log n})$ (see Definition 14). Then $a_{1},\ldots,a_{n}$ satisfy the follows:
@@
    (Bounded maximum norm) With probability at least $1-\binom{n}{d}^{-1}$, $\max_{i\in n}\|a_{i}-\bar{a}_{i}\|\leq 4\sigma\sqrt{d\log n}\cdot\max_{i\in[n]}\|\bar{a}_{i}\|$.
@@
    (Bounded expected radius of projection) For any $k\leq d$, any fixed $k$-dimensional linear subspace $H\subseteq\mathbb{R}^{d}$, we have $\mathbb{E}[\max_{i\in[n]}\|\pi_{H}(a_{i}-\bar{a}_{i})\|]\leq 4\sigma\sqrt{k\log n}\cdot\max_{i\in[n]}\|\bar{a}_{i}\|$.
@@
Although [DH18] state the above lemma only for $d\geq 3$, their proof applies without change to the case $d=2$.
@@ 2.3 Change of Variables
We will use the following change of variables, which is a standard tool in stochastic and integral geometry.
Definition 18 (Change of variables).
Let $a_{1},\ldots,a_{d}$ be $d$ affine independent vectors in $\mathbb{R}^{d}$. Let $\theta\in\mathbb{S}^{d-1},t\in\mathbb{R}$ be such that $\forall i\in[d],\theta^{\top}a_{i}=t$ and suppose $\theta^{\top}e_{1}>1$ without loss of generality where $e_{1}=(1,0,\ldots,0)^{\top}\in\mathbb{R}^{d}$ is the unit vector that has nonzero element on the first coordinate.
@@
$\phi(\theta,t,b_{1},\ldots,b_{d})=(R_{\theta}(b_{1})+t\theta,\ldots,R_{\theta}(b_{d})+t\theta)=(a_{1},\ldots,a_{d}).$ (12)
Lemma 19 (Jacobian of the inverse isometric transformation, see Theorem 7.2.7 in [Schneider2008]).
Let $\phi:\mathbb{S}^{d-1}\times\mathbb{R}\times\mathbb{R}^{(d-1)\times d}\to\mathbb{R}^{d\times d}$ be the transformation defined in Definition 18. The inverse transformation of $\phi$ is defined almost everywhere and has Jacobian determinant that equals to
$\left|\det\left(\frac{\partial\phi(a)}{\phi(a)}\right)\right|=C_{d}\cdot(d-1)!\cdot\operatorname{vol}_{d-1}(\operatorname{conv}(a_{1},\ldots,a_{d}))$
for some constant $C_{d}$ depending only on the dimension. As a consequence, if $a_{1},\dots,a_{d}$ are points with probability density $\mu(a_{1},\dots,a_{d})$ and if $\theta\in\mathbb{S}^{d-1},t\in\mathbb{R},b_{1},\dots,b_{d}\in\mathbb{R}^{d-1}$ have probability density proportional to
@@
Lemma 20 (Log-Lipschitzness of the Position of Affine Hull).
Let $a_{1},\ldots,a_{d}\in\mathbb{R}^{d}$ be $d$ independent $L$-log-Lipschitz random variables, and let $(\theta,t,b_{1},\dots,b_{d})=\phi^{-1}(a_{1},\dots,a_{d})$, where $\phi:\mathbb{S}^{d-1}\times\mathbb{R}\times\mathbb{R}^{(d-1)\times d}\to\mathbb{R}^{d\times d}$ is defined in Definition 18. Then conditional on the values of $\theta,b_{1},\dots,b_{d}$, the random variable $t$ is $(dL)$-log-Lipchitz.
Proof.
By Lemma 19, the joint probability density of $(a_{1},\ldots,a_{d})$ is proportional to
@@
where $\mu_{i}$ is the probability density of $a_{i}$. Conditioning on $b_{1},\dots,b_{d}\in\mathbb{R}^{d-1}$ and $\theta\in\mathbb{S}^{d-1}$, the volume $\operatorname{vol}_{d-1}(\operatorname{conv}(b_{1},\ldots,b_{d}))$ is fixed. The statement then follows from the fact that for each $i\in[d]$, $\mu_{i}(R_{\theta}(b_{i})+t\theta)$ is $L$-log-Lipschitz in $t$ for any $b_{i}$. ∎
@@ 2.4 Non-Degenerate Conditions
Definition 21 (Non-degenerate polytope).
@@
Definition 22 (Non-degenerate intersection with a 2D-plane).
@@
    the vertices of $Q\cap W$ have one-to-one correspondence to the $(d-2)$-dimensional faces (ridges) of $Q$ that have non-empty intersection with $W$
Fact 23 (Non-degenerate conditions of random polytope).
Given any $n\geq d\geq 2$ and any fixed two-dimensional plane $W\subseteq\mathbb{R}^{d}$. For $a_{1},\ldots,a_{n}\in\mathbb{R}^{d}$, the polytope $Q=\operatorname{conv}(a_{1},\ldots,a_{n})$ satisfies the following properties everywhere except for a set of measure 00:
@@
Assume the polytope $Q=\operatorname{conv}(a_{1},\dots,a_{n})$ and the two-dimensional linear subspace $W\subseteq\mathbb{R}^{d}$ satisfy the non-degenerate conditions in 23. Each edge of the two-dimensional polygon formed by the intersection $W\cap Q$ can be described by a set of $d$ vertices, where the edge is equivalent to the intersection of $W$ with the convex hull of these $d$ vertices. Furthermore, each vertex of $W\cap W$ corresponds to a set of $(d-1)$ vertices. The following lemma characterizes the relation of these sets for adjacent vertices and edges:
Fact 24 (Properties of neighboring vertices on non-degenerate intersection polygon).
@@
Next we consider $|I\backslash J_{2}|$. Since $J_{1}\subseteq I$ and $|J_{1}\backslash J_{2}|=1$, it could only be the case that $|I\backslash J_{2}|\in\{1,2\}$. If $|I\backslash J_{2}|=1$, then by $|I|=|J_{2}|+1$ we must have $J_{2}\subseteq I$, but this contradicts to the fact that $\operatorname{conv}(J_{2})\not\subseteq\operatorname{conv}(I)$. Therefore we could only have $|I\backslash J_{2}|=2$. ∎

### v3

@@ 1.3.2 Smoothed Complexity Lower Bound
Our smoothed complexity lower bound (Theorem 6.1) is based on two geometric observations using the inner and outer radius of the perturbed polytope. For a polytope $P$ and a unit norm ball $\mathbb{B}$, its outer radius with center $x$ is the smallest $R$ such that there exists $P\subseteq R\cdot\mathbb{B}+x$. Its inner radius with center $x$ is the largest $r$ such that $r\cdot\mathbb{B}+x\subseteq P$.
The first observation is that, if a two-dimensional polygon $T$ has inner $\ell_{2}$-radius of $r$ and outer $\ell_{2}$-radius of $(1+\varepsilon)\cdot r$ with respect to the same center, then $T$ has at least $\Omega(\varepsilon^{-1/2})$ edges (Lemma 6.16). This comes from the fact that every edge of $T$ has length at most $O(r\sqrt{\varepsilon})$, whereas the perimeter of $T$ is at least $2\pi r$.
Second, if two polytopes $Q,\tilde{Q}\subseteq\mathbb{R}^{d}$, each with inner radius $t$, have Hausdorff distance $\varepsilon<t/2$ to each other, then $Q$ will approximate $\tilde{Q}$ in the way that (Lemma 6.14)
@@
To prove our lower bound, we construct a polytope $Q=\operatorname{conv}(\bar{a}_{1},\ldots,\bar{a}_{n})\subseteq\mathbb{R}^{d}$ and a two-dimensional linear subspace $W$ such that $\Omega(1)\cdot\mathbb{B}^{d}_{1}\subseteq Q\subseteq\mathbb{B}^{d}_{1}$, and $Q\cap W$ has outer $\ell_{2}$-radius $r>0$ and inner $\ell_{2}$-radius $\frac{r}{(1+4^{-d})}$. Perturbing the vertices of $Q$, we obtain $\tilde{Q}=\operatorname{conv}(a_{1},\ldots,a_{n})$, where $a_{i}\sim N(\bar{a}_{i},\sigma^{2}I_{d\times d})$ for each $i\in[n]$. Note that $Q\subseteq\mathbb{B}_{1}^{d}$ implies that $\bar{a}_{1},\dots,\bar{a}_{n}$ satisfy the normalization requirement in (1). With high probability the Hausdorff distance in $\ell_{1}$ between $Q$ and $\tilde{Q}$ is bounded by $\max_{i\in[n]}\|a_{i}-\tilde{a}_{i}\|_{1}\leq O(\sigma d\sqrt{\log n})$. Using (10), we bound the inner and outer radius of $\tilde{Q}\cap W$. A lower bound on the number of edges of $\tilde{Q}\cap W$ then follows from Lemma 6.16 as described above.
We remark that the polytopes $Q=\operatorname{conv}(a_{1},\dots,a_{n})\subseteq\mathbb{R}^{d}$ with $n=O(d)$ and two-dimensional subspaces $W$ such that $Q\cap W$ has inner $\ell_{2}$-radius $\frac{r}{1+4^{-d}}$ and outer $\ell_{2}$-radius $r>0$ were first obtained by [8] as an extended formulation for a regular $2^{k}$-gon with $O(k)$ variables and $O(k)$ inequalities. Their polytope, however, has an outer and inner radius that differ by a factor $2^{\Omega(k)}$, meaning that we cannot apply Lemma 6.14 for $\sigma>2^{-k}$. We construct an alternative extended formulation where the ratio between inner and outer $\ell_{1}$-radius is only $O(1)$. With an appropriate scaling to get $Q\subseteq B_{1}^{d}$, we find that the perturbed polytope $\tilde{Q}$ will have intersection $\tilde{Q}\cap W$ with inner radius $\frac{r}{1+4^{-d}}(1-2\varepsilon/t)$ and outer radius $(1+\varepsilon/t)r$, where $\varepsilon=O(\sigma d\sqrt{\log n})$, and thus has $\Omega(\min(\frac{1}{\sqrt{\varepsilon}},2^{d}))$ edges, with high probability.
@@ 2 Preliminaries
We write $n]:=\{1,\dots,n\}$. By $\operatorname{conv}(a_{1},\dots,a_{n})=\operatorname{conv}(a_{i}:i\in[n])$ we denote the convex hull of vectors $a_{1},\dots,a_{n}$, and similarly we write for the affine hull as $\operatorname{affhull}(a_{i}:i\in[n])$. For sets $A,B\subseteq\mathbb{R}^{d}$, the distance between the two is $\operatorname{dist}(A,B)=\inf_{a\in A,b\in B}\|a-b\|$. For a point $x\in\mathbb{R}^{d}$ we write $\operatorname{dist}(x,A)=\operatorname{dist}(A,x)=\operatorname{dist}(A,\{x\})$.
@@
For a convex body $K\in\mathbb{R}^{d}$, we define $\partial K\subseteq\mathrm{span}(K)$ as the boundary of $K$ in the linear subspace spanned by the vectors in $K$.
@@ 2.1 Polytopes
Definition 2.1 (Polytope).
A convex set $P\subseteq\mathbb{R}^{d}$ is a polyhedron if it can be expressed as $P=\{x\in\mathbb{R}^{d}:Ax\leq b\}$ for some $A\in\mathbb{R}^{n\times d}$, $b\in\mathbb{R}^{n}$ where $n\in\mathbb{Z}_{+}$. A bounded polyhedron is also called a polytope.
Definition 2.2 (Valid Condition and Facet).
@@
A subset $F\subseteq P$ is called a face of $P$ if $F=P\cap\{x\in\mathbb{R}^{d}:x^{\top}c=d\}\neq\emptyset$ for some valid condition $x^{\top}c\leq d$. A facet is a d1d-1-dimensional face, a ridge is a d2d-2-dimensional face, an edge is a $1$-dimensional face and a vertex is a 0-dimensional face.
Definition 2.3 (Polar dual of a convex body).
@@
Fact 1 (Polar dual of polytope).
@@ 2.2 Probability Distributions
Definition 2.4 (Gaussian distribution).
The $d$-dimensional Gaussian distribution $\mathcal{N}_{d}(\bar{a},\sigma^{2}I)$ with support $\mathbb{R}^{d}$, mean $\bar{a}\in\mathbb{R}^{d}$, and standard deviation $\sigma$ is defined by the probability density function
@@
Lemma 2.5 (Gaussian tail bound).
@@
Corollary 2.6 (Global diameter of Gaussian random variables).
@@
Proof 2.7.
From Lemma 2.5, we have for each $i\in[n]$ that
@@
Then the statement follows from the union bound over all choices of $i\in[n]$.
A helpful technical substitute for the Gaussian distribution was introduced by [17]:
Definition 2.8 ($(\sigma,r)$-Laplace-Gaussian distribution).
@@
Definition 2.9 ($L$-log-Lipschitz random variable).
@@
Lemma 2.10 (Properties of Laplace-Gaussian random variables, Lemmas 3.7 and 3.33 of [17]).
Given any $n\geq d$, $\sigma>0$. Let $a_{1},\ldots,a_{n}\in\mathbb{R}^{d}$ be independent random variables each sampled form $LG_{d}(\bar{a},\sigma,4\sigma\sqrt{d\log n})$ (see Definition 2.8). Then $a_{1},\ldots,a_{n}$ satisfy the follows:
@@
    (Bounded maximum norm) With probability at least $1-\binom{n}{d}^{-1}$, $\max_{i\in n}\|a_{i}-\bar{a}_{i}\|\leq 4\sigma\sqrt{d\log n}$.
@@
    (Bounded expected radius of projection) For any $k\leq d$, any fixed $k$-dimensional linear subspace $H\subseteq\mathbb{R}^{d}$, we have $\mathbb{E}[\max_{i\in[n]}\|\pi_{H}(a_{i}-\bar{a}_{i})\|]\leq 8\sigma\sqrt{k\log n}$.
@@
Although [17] state the above lemma only for $d\geq 3$, their proof applies without change to the case $d=2$.
@@ 2.3 Change of Variables
We will make use of a specific change of variables, which is a standard tool in stochastic and integral geometry. It will allow us to investigate, for points $a_{1},\dots,a_{d}\in\mathbb{R}^{d}$, the distribution of these points when conditioning on a specific affine hyperplane $\operatorname{affhull}(a_{1},\dots,a_{d})$. This will enable us to make conclusions about the shapes of the faces of the convex hull $\operatorname{conv}(a_{1},\dots,a_{n})$ of $n$ vectors.
Definition 2.12 (Change of variables).
Let $\theta\in\mathbb{S}^{d-1},t\in\mathbb{R}$ be such that $\forall i\in[d],\theta^{\top}a_{i}=t$ and suppose $\theta^{\top}e_{1}>1$ without loss of generality where $e_{1}=(1,0,\ldots,0)^{\top}\in\mathbb{R}^{d}$ is the unit vector that has nonzero element on the first coordinate.
@@
$\phi(\theta,t,b_{1},\ldots,b_{d})=(R_{\theta}(b_{1})+t\theta,\ldots,R_{\theta}(b_{d})+t\theta).$ (12)
The choice of the embeddings $\tilde{R}_{\theta}$ is largely arbitrary. What matters is that the transformation $\phi$ and its inverse are continuous almost everywhere.
Lemma 2.13 (Jacobian of the transformation, see Theorem 7.2.7 in [46]).
Let $\phi:\mathbb{S}^{d-1}\times\mathbb{R}\times\mathbb{R}^{(d-1)\times d}\to\mathbb{R}^{d\times d}$ be the transformation defined in Definition 2.12. The transformation $\phi$ is defined almost everywhere and has Jacobian determinant that equals to
$\left|\det\left(\frac{\partial\phi}{\partial(\theta,t,b_{1},\dots,b_{d})}\right)\right|=C_{d}\cdot(d-1)!\cdot\operatorname{vol}_{d-1}(\operatorname{conv}(b_{1},\ldots,b_{d}))$
for some constant $C_{d}$ depending only on the dimension. As a consequence, if $a_{1},\dots,a_{d}\in\mathbb{R}^{d}$ are points with probability density $\mu(a_{1},\dots,a_{d})$ and if $\theta\in\mathbb{S}^{d-1},t\in\mathbb{R},b_{1},\dots,b_{d}\in\mathbb{R}^{d-1}$ have probability density proportional to
@@
The interested reader might observe that if $(a_{1},\dots,a_{d})=\phi(\theta,t,b_{1},\dots,b_{d})$ then the Jacobian
$(d-1)!\cdot\operatorname{vol}_{d-1}(\operatorname{conv}(a_{1},\ldots,a_{d}))=(d-1)!\cdot\operatorname{vol}_{d-1}(\operatorname{conv}(b_{1},\ldots,b_{d}))$
is equal to the determinant of the $d\times d$ matrix $(\theta^{\top},(a_{1}-a_{d})^{\top},(a_{2}-a_{d})^{\top},\dots,(a_{d-1}-a_{d})^{\top})$, which is equal to the determinant of the $(d-1)\times(d-1)$ matrix with rows or columns given by $b_{1}-b_{d},b_{2}-b_{d},\dots,b_{d-1}-b_{d}$.
@@
Lemma 2.14 (Log-Lipschitzness of the Position of Affine Hull).
Let $a_{1},\ldots,a_{d}\in\mathbb{R}^{d}$ be $d$ independent $L$-log-Lipschitz random variables, and let $(\theta,t,b_{1},\dots,b_{d})=\phi^{-1}(a_{1},\dots,a_{d})$, where $\phi:\mathbb{S}^{d-1}\times\mathbb{R}\times\mathbb{R}^{(d-1)\times d}\to\mathbb{R}^{d\times d}$ is defined in Definition 2.12. Then conditional on the values of $\theta,b_{1},\dots,b_{d}$, the random variable $t$ is $(dL)$-log-Lipchitz.
Proof 2.15.
By Lemma 2.13, the joint probability density of $(a_{1},\ldots,a_{d})$ is proportional to
@@
where $\mu_{i}$ is the probability density of $a_{i}$. Conditioning on $b_{1},\dots,b_{d}\in\mathbb{R}^{d-1}$ and $\theta\in\mathbb{S}^{d-1}$, the volume $\operatorname{vol}_{d-1}(\operatorname{conv}(b_{1},\ldots,b_{d}))$ is fixed. The statement then follows from the fact that for each $i\in[d]$, $\mu_{i}(R_{\theta}(b_{i})+t\theta)$ is $L$-log-Lipschitz in $t$ for any $b_{i}$.
Conditional on $\theta$ and $t$, the points $b_{1},\dots,b_{d}$ are not $L$-log-Lipschitz due to the volume term. When it becomes relevant, we will show that this factor does not affect the argument in any negative way.
@@ 2.4 Non-Degenerate Conditions
Definition 2.16 (Non-degenerate polytope).
@@
Definition 2.17 (Non-degenerate intersection with a 2D-plane).
@@
    the vertices of $Q\cap W$ have one-to-one correspondence to the (d2)(d-2)-dimensional faces (ridges) of $Q$ that have non-empty intersection with $W$
Fact 4 (Non-degenerate conditions of random polytope).
Given any $n\geq d\geq 2$ and any fixed two-dimensional plane $W\subseteq\mathbb{R}^{d}$. For $a_{1},\ldots,a_{n}\in\mathbb{R}^{d}$, the polytope $Q=\operatorname{conv}(a_{1},\ldots,a_{n})$ satisfies the following properties everywhere except for a set of measure 0:
@@
Assume the polytope $Q=\operatorname{conv}(a_{1},\dots,a_{n})$ and the two-dimensional linear subspace $W\subseteq\mathbb{R}^{d}$ satisfy the non-degenerate conditions in Fact 4. Each edge of the two-dimensional polygon formed by the intersection $W\cap Q$ can be described by a set of $d$ vertices, where the edge is equivalent to the intersection of $W$ with the convex hull of these $d$ vertices. Furthermore, each vertex of $Q\cap W$ corresponds to a set of (d1)(d-1) vertices. The following lemma characterizes the relation of these sets for adjacent vertices and edges:
Fact 5 (Properties of neighboring vertices on non-degenerate intersection polygon).
@@
Next we consider $|I\backslash J_{2}|$. Since $J_{1}\subseteq I$ and $|J_{1}\backslash J_{2}|=1$, it could only be the case that $|I\backslash J_{2}|\in\{1,2\}$. If $|I\backslash J_{2}|=1$, then since $|I|=|J_{2}|+1$ we must have $J_{2}\subseteq I$, but this contradicts to the fact that $\operatorname{conv}(J_{2})\not\subseteq\operatorname{conv}(I)$. Therefore we could only have $|I\backslash J_{2}|=2$.